# Swimming mode determines how well mesoscale swimmers shield their odor in turbulence

### v1

@@ Abstract
Marine organisms manipulate their surrounding flow through their swimming dynamics, which affects the transport of their own odor cues. We demonstrate by direct numerical simulations how a group of mesoscale swimmers immersed in a turbulent flow alters the shape of the odor plume they release in the water. Odor mixing is enhanced by increased velocity fluctuations and a swimmer-induced flow circulation which widen the odor plume at close range while speeding up dilution of the chemical trace. Beyond a short-range increase in the likelihood of being detected, swimming considerably reduces detections with effects that can persist at distances of the order of ten times the size of the group or more. We find that puller-like swimmers are more effective at olfactory shielding than pusher-like swimmers. We trace this difference back to the dynamics at the swimmer location, which tends to trap odor at the source for pushers and to dilute it for pullers. Olfactory shielding is robust to changes in the conditions, and is more pronounced for weak turbulent Reynolds numbers and large swimmer Reynolds numbers. Our results suggest that olfactory shielding may play a role in the emergence of different swimming modalities by marine organisms.
@@ I Introduction
Figure 1: A snapshot of the odor field emitted by the swimmers. The swimmers are placed in a cuboidal box of dimension $100\eta\times 100\eta\times 25\eta$ near the inlet, as shown in the inset. The simulation box has a dimension of $8000\eta\times 1000\eta\times 400\eta$.
@@
Groups of swimming organisms are a major target for large predators who employ multiple sensory cues to track them down. In particular, swimmers give out their whereabouts by shedding odor, which is carried downstream of the group by water flow and can be detected by predators even from large distances (see, e.g., [1, 2] and references therein). As odor is transported away from the group, it undergoes turbulent mixing, which merges odor-laden and odor-free water. Mixing will ultimately erase their chemical trace as odor gets diluted below the threshold of detection of potential predators. The act of swimming in itself enhances turbulent mixing by actively manipulating the surrounding flow in a process known as biomixing, as documented by laboratory and numerical studies (see, e.g., [3] and references therein). Biomixing may even affect the global ocean circulation, although its quantification is challenging and hotly debated [4, 5]. As a direct consequence of the documented enhanced mixing one expects that odor will dilute more efficiently, speeding up the elimination of the swimmers’ own odor trace, as previously hypothesized for filter feeding by sessile organisms [6, 7]. Given the fascinating diversity of species in the ocean, one naturally wonders whether specific ways of swimming may be more efficient than others at this mechanism, that we refer to as olfactory shielding.
@@
Swimming organisms leverage a diversity of self-propulsion mechanisms, e.g. beating or rotating appendages like cilia or flagella or propagating shape deformations [8, 9]. Mesoscale swimmers such as copepods, amphipods and other small crustaceans form an important part of the marine ecosystem. They serve as pivotal food sources for larger marine species and play an important role in marine ecology [10]. Such species are often found in large groups [11] and inhabit regions of mild-to-moderate turbulence [12, 13, 8]. At the micro- and meso-scales, swimmers can exhibit pusher or puller dynamics, depending on whether they generate thrust at their rear or front during specific swimming modes [14, 15]. The hydrodynamic interactions by pushers and pullers generate distinct patterns in their surrounding water flow, both individually [16, 17, 15] as well as in groups [18, 19]. But whether and how mesoscale pushers vs pullers are more or less efficient at olfactory shielding remains to be understood.
@@
In this article, we provide a detailed analysis of the coupling of mechanical and chemical signals from mesoscale swimmers in turbulent flows. We conduct high-resolution direct numerical simulations using a model system of a collection of point swimmers in an open channel turbulent flow. We show that groups of swimmers enhance mixing, and puller-like swimmers are more efficient than pusher-like swimmers at olfactory shielding. We demonstrate that this effect is stronger for large swimmer Reynolds numbers and low flow Reynolds numbers, suggesting that olfactory shielding may contribute selective pressure on mesoscale swimmers in the ocean.
@@ II Modeling and methodology
We consider a collection of mesoscale swimmers moving at a uniform, constant velocity, embedded in an open channel flow. The fluid flow is modeled through direct numerical simulation of the incompressible Navier-Stokes equation
@@
$\rho\left(\frac{\partial\bm{u}}{\partial t}+\bm{u}\cdot\nabla\bm{u}\right)$ $=-\nabla P+\mu\nabla^{2}\bm{u}+\bm{f}+\bm{f}_{p};$ (1)
$\nabla\cdot\bm{u}$ $=0.$
Here, $\bm{u}$ is the fluid velocity field, $P$ is the pressure, $\rho$ is the fluid density, $\mu$ is the dynamic viscosity and $\bm{f}_{p}$ is the forcing due to the swimmer dynamics. We solve equation (1) using a second-order central finite-difference scheme on a staggered grid [20].
Figure 2: Statistics of the mean odor field. The contour plots show how the range of the mean odor field varies between pusher-like, puller-like and neutral swimmers for vertical and horizontal mid-plane cross-sections. (a) Vertical cross-sections of the mean odor differential $\Delta\bar{c}(\bm{x})$ for pusher-like swimmers and puller-like swimmers and the corresponding odor standard deviation differential $\Delta\sigma(\bm{x})$. The region containing the swimmers is marked by a red box. (b) The corresponding figures for the horizontal cross-section. (c) Odor profiles at different distances from the swimmers’ location. (d) Shielding intensities of the mean odor field $\Delta V(C)/V_{0}(C)$. (e) Shielding intensity of the odor fluctuations $\Delta V^{\sigma}(s)/V^{\sigma}_{0}(s)$ for pusher-like and puller-like swimmers, showing that the swimming dynamics dampens large fluctuations.
The channel is forced with a constant pressure gradient $\bm{f}$ [21]. The boundary conditions for the channel flow are as follows: a fixed velocity on the bottom boundary (to represent the appropriate swimming velocity of the swimmers, as the simulations are carried out in the reference frame of the swimmers), free slip on the top boundary ($u_{z}=0,\partial_{z}u_{x}=\partial_{z}u_{y}=0$) and periodic in the other directions. We simulate three different fluid friction Reynolds numbers $Re_{\tau}=u_{\tau}H/\nu$, where $u_{\tau}$ is the shear velocity, $H$ is the channel height and $\nu$ is the kinematic viscosity ($\nu=\mu/\rho$), as shown in Table 1. We choose Environment 1 in Table 1 as our base environment and all units are normalized by using the Kolmogorov scales in this simulation. The simulation box has a dimension of $8000\eta\times 1000\eta\times 400\eta$.
@@
$\frac{\partial c}{\partial t}+\bm{u}\cdot\nabla c=D\nabla^{2}c+s.$ (2)
Here $c$ is odor concentration, $D$ is the diffusivity and $s$ is the odor source. We assume perfect adsorption on the bottom boundary ($c=0$), no flux at the top boundary ($\partial_{z}c=0$) and outflow conditions at the other boundaries. We use a constant source term at the position of the swimmers. For our simulations, we use Schmidt number $Sc=1$. Although $Sc$ in water is typically higher than this, we anticipate a weak dependence on $Sc$ since we are interested in large-scale statistics [22, 23].
Figure 3: The effect of swimming dynamics on the detection probability. (a) $\Delta P(\bm{x};C_{H})$ for puller-like swimmers. (b) and (c) Odor pdf differential $\Delta f(c)$ in regions A and B, respectively, for puller-like swimmers. (d) Odor pdf differential $\Delta f(c)$ for both types of swimmers; statistics in regions A and B are represented with solid and dotted lines respectively. (e) Detection probability $P_{x}(x;C_{H})$ as a function of stream-wise distance, averaged along $y$ and $z$. The $x-$axis starts at $x=500$, where the group of swimmers ends. At long distances $x\gtrsim 700\,\eta$ the detection probability is reduced due to the swimming dynamics. Shielding persists up to $x\sim 2500\,\eta$
To model the forcing $\bm{f}_{p}$ due to the swimmers, note that the swimmers move at a constant velocity thus the net force on them is zero. The flow field due to neutrally buoyant micro and mesoscale swimmers can be approximated by a force dipole [24, 25, 26]. Further addition of torque dipoles combined with rigid boundary conditions can effectively capture swimmers moving at even higher $Re$ [27]. However, since our swimmers are of the order of the Kolmogorov scale and move at relatively low $Re$, we model our swimmers using a force dipole approximation. Each swimmer is modeled as a force dipole oriented along the streamwise direction $F(\delta(\bm{x}+\bm{r})-\delta(\bm{x}-\bm{r}))\hat{\bm{x}}$, where $r$ is half the length of the swimmer and $F$ is the magnitude of the forcing. The orientation of the force dipole determines whether the swimmer is pusher-like or puller-like. When the forcing points outwards relative to the swimmer’s axis, the swimmer “pushes” the fluid away and is a pusher-like swimmer whereas, when it points inwards, it “pulls” the fluid in, thus modeling a puller-like swimmer [15]. The magnitude of $F$ is computed based on the experimental results for the drag force at intermediate Reynolds numbers [28] (also see Eq. S1 in the Supporting Information). The swimmers are placed in a cuboidal box of dimension $100\eta\times 100\eta\times 25\eta$ close to the inlet with a total of $10\times 20\times 5$ swimmers along the streamwise, spanwise and wall-normal directions respectively.
@@
The list of environments is given in Table 1. All figures correspond to the base environment (Simulation 1 in Table 1) unless otherwise mentioned. For each of these environments, we conduct three distinct simulations: In the first two, we incorporate the dynamics of the swimmers by accounting for the forcing imposed by the swimmers on the fluid – pusher-like or puller-like; then we remove this forcing, providing a baseline against which to measure the swimmers’ effects. The simulation setup is shown in Fig. 1, illustrating a typical case of odor dispersion in a channel flow.
@@ III Results
To determine whether swimmers’ dynamics leave their signatures in the odor distribution over long time-scales, we begin our analysis by examining the effect of hydrodynamic interactions on the mean and standard deviation of the odor field (Fig. 2). Since the odor distribution is inhomogeneous and statistically stationary, we approximate the ensemble average of the mean and standard deviation of the odor field with temporal averages:
$\bar{c}(\bm{x})$ $=\langle c(\bm{x},t)\rangle$ (3)
$\sigma(\bm{x})$ $=\langle(c(\bm{x},t)-\bar{c}(\bm{x},t))^{2}\rangle^{\frac{1}{2}}.$
To account for the effect of the swimmers, we now define differential quantities by comparison with the baseline simulation that ignores the presence of the swimmers:
$\Delta\bar{c}_{s}(\bm{x})$ $=[\bar{c}_{s}(\bm{x})-\bar{c}_{0}(\bm{x})]/C_{\Omega}$ (4)
$\Delta\sigma_{s}(\bm{x})$ $=[\sigma_{s}(\bm{x})-\sigma_{0}(\bm{x})]/C_{\Omega},$
where the subscript $s$ stands for swimmers (either pusher-like or puller-like), and the subscript 0 represents the baseline simulation. Mean and fluctuations are normalized by $C_{\Omega}=\text{max}(\bar{c}_{0})$, which occurs within the volume $\Omega$ containing the swimmers.
@@
To quantify the overall effect of swimmers in the entire volume, we define the volumes $V(C)$ where the mean odor concentration exceeds $C$ and $V^{\sigma}(s)$ where the fluctuations exceed $s$:
$V_{s}(C)=\int H(\bar{c}_{s}(\bm{x})-C)\text{d}{\bm{x}},$
$V_{s}^{\sigma}(s)=\int H(\sigma_{s}(\bm{x})-s)\text{d}{\bm{x}},$
where $H(x)$ is the Heaviside step function and integration is extended to the entire domain. We define the shielding efficiency as the decrease in the volumes defined above:
$\Delta V_{s}(C)=V_{s}(C)-V_{0}(C),$
$\Delta V_{s}^{\sigma}(s)=V^{\sigma}_{s}(C)-V^{\sigma}_{0}(s).$
When $C\approx 0$, swimming dynamics has little to no effect (or even a weak negative effect) in shielding the mean odor, with puller-like swimmers performing slightly worse than pusher-like ones (Fig. 2 (d)). As $C$ is increased, $\Delta V(C)<0$ thus both types of swimmers decrease the total volume where the average concentration is above the threshold. For a threshold of $C/C_{\Omega}=0.5$, pullers shrinks $V(C)$ by $50\%$, while pushers shrink $V(C)$ by $45\%$. Similarly, shielding efficiency for odor fluctuations (Fig. 2 (e)) shows that regions with extreme fluctuations are significantly reduced when the swimmer dynamics are taken into account. Interestingly, puller-like swimmers exhibit better shielding of both mean and fluctuations, which may reduce the chances of detection.
Figure 5: (a) Shielding of odor mean $\Delta V(C)/V_{0}(C)$ as a function of normalized threshold $C/C_{\Omega}$, (b) shielding of odor fluctuations $\Delta V^{\sigma}(s)/V^{\sigma}_{0}(s)$ as a function of normalized threshold $s/C_{\Omega}$, (c) normalized detection probability differential $\Delta P_{x}(x;C_{H})/P_{x}0(x;C_{H})$ as a function of the streamwise distance from the swimmers and (d) odor pdf differential $\Delta f(c;\Omega_{A}/B)$. Colors represent different swimmer Reynolds numbers (a-d) or different turbulent Reynolds numbers (e-h). Pusher-like and puller-like swimmers are represented by solid and dotted lines respectively.
@@
Unintuitively, the two swimmer types have opposite effects at the odor source, yet they both shrink their own detection probability at large enough distances from the group. This is best summarized by monitoring the probability of detection as a function of distance $x$ downstream of the swimmers:
$P_{x}(x,C)$ $=\frac{1}{L_{z}L_{y}}\int_{0}^{L_{z}}\int_{0}^{L_{y}}P(\bm{x},C)\text{d}y\text{d}z.$ (7)
The group increases their probability of detection at close range, i.e. within $\sim$ $200\eta$ from the end of the group, about twice the length of the group. At longer range the group decrease their detection probability (Fig. 3 (e)). Importantly, shielding persists up to $\sim$ $2000\eta$, i.e. 20 times the length of the group. To reconcile this unintuitive mismatch, we analyze the full probability density function within regions A and B and compare the simulations with and without swimmers $\Delta f_{s}(c,\bm{x})=f_{s}(c,\bm{x})-f_{0}(c,\bm{x})$. In region A the swimmer dynamics dampen both extremes of the odor distribution (Fig. 3 (b)). This is a signature of increased mixing, which homogenizes the odor field towards its mean concentration. In region B the swimmer dynamics dilutes the odor, presumably by spreading it in a larger region further outward (Fig. 3 (c)). Pusher-like swimmers behave qualitatively similarly (Fig. 3 (d)). These signatures of enhanced mixing suggest that the velocity fluctuations due to the swimmer dynamics play a major role in odor transport.

### v2

# Effect of swimming mode on shielding of odor traces in turbulence

@@ Abstract
Marine organisms manipulate their surrounding flow through their swimming dynamics, which affects the transport of their own odor cues. We demonstrate by direct numerical simulations how a group of swimmers, moving at intermediate Reynolds numbers, immersed in a turbulent flow, alter the shape of the odor plume they release in the water. Odor mixing is enhanced by increased velocity fluctuations and a swimmer-induced flow circulation which widens the odor plume at close range while speeding up dilution of the chemical trace. Beyond a short-range increase in the likelihood of being detected, swimming considerably reduces detections with effects that can persist at distances of the order of ten times the size of the group or more. We find that puller-like swimmers are more effective at olfactory shielding than pusher-like swimmers. We trace this difference back to the dynamics at the swimmer location, which tends to trap odor at the source for pushers and to dilute it for pullers. Olfactory shielding is robust to changes in the conditions, and is more pronounced for weak turbulent Reynolds numbers and large swimmer Reynolds numbers. Our results suggest that olfactory shielding may play a role in the emergence of different swimming modalities by marine organisms.
@@ I Introduction
Figure 1: A snapshot of the odor field emitted by the swimmers. The swimmers are placed in a cuboidal box of dimension $100\eta\times 100\eta\times 25\eta$ near the inlet, as shown in the inset (here $\eta$ is the Kolmogorov length scale). The simulation box has a dimension of $8000\eta\times 1000\eta\times 400\eta$.
@@
Groups of swimming organisms are a major target for large predators who employ multiple sensory cues to track them down. In particular, swimmers give out their whereabouts by shedding odor, which is carried downstream of the group by water flow and can be detected by predators even from large distances (see, e.g., [43, 30] and references therein). As odor is transported away from the group, it undergoes turbulent mixing, which merges odor-laden and odor-free water. Mixing will ultimately erase their chemical trace as odor gets diluted below the threshold of detection of potential predators. The act of swimming in itself enhances turbulent mixing by actively manipulating the surrounding flow in a process known as biomixing, as documented by laboratory and numerical studies (see, e.g., [21] and references therein). Biomixing may even affect the global ocean circulation, although its quantification is challenging and hotly debated [18, 40]. As a direct consequence of the documented enhanced mixing one expects that odor will dilute more efficiently, speeding up the elimination of the swimmers’ own odor trace, as previously hypothesized for filter feeding by sessile organisms [9, 2]. Given the fascinating diversity of species in the ocean, one naturally wonders whether specific ways of swimming may be more efficient than others at diluting their odor trace, which we refer to as olfactory shielding. Evidence that swimming kinematics can indeed shape chemical cue transport already exists; for example, studies have shown that undulatory swimming patterns can enhance the dispersal of odor cues [17].
Swimming organisms leverage a diversity of self-propulsion mechanisms, e.g. beating or rotating appendages like cilia or flagella or propagating shape deformations [14, 44]. Swimmers such as copepods, amphipods and other small crustaceans which move at intermediate Reynolds number ($Re$) form an important part of the marine ecosystem [20]. In this manuscript we refer to them as “mesoscale” swimmers: note that this use of the term mesoscale is unrelated to its use in oceanography, where mescoscales are intermediate spatial scales. Mesoscale swimmers serve as pivotal food sources for larger marine species and play an important role in marine ecology [29]. Such species are often found in large groups [12], inhabit regions of mild-to-moderate turbulence [23, 26, 14] and show a fascinating variety in behavior [4, 5, 24, 27]. Field and laboratory observations suggest that several species of crustaceans form dense swarms composed of thousands of individuals that can maintain their position in mild turbulence [7, 6, 11, 35]. For instance, Dioithona oculata swarms withstand currents up to 2 cm/s (fluid Re>4000Re>4000[7, 6, 3], while several mysid species aggregate in shallow surge channels during day and night [11, 25].
@@
Swimmers moving at low and intermediate Reynolds number can exhibit pusher or puller dynamics, depending on whether they generate thrust at their rear or front during specific swimming modes [34, 22]. The hydrodynamic interactions by pushers and pullers generate distinct patterns in their surrounding water flow, both individually [19, 36, 22] as well as in groups [45, 8]. But whether and how mesoscale pushers vs pullers are more or less efficient at olfactory shielding remains to be understood.
@@
In this article, we provide a detailed analysis of the coupling of mechanical and chemical signals from mesoscale swimmers in shallow, turbulent flows, drawing inspiration from the collective dynamics of crustacean swarms [7, 6, 11]. We conduct high-resolution direct numerical simulations using a model system of a collection of point swimmers in an open channel turbulent flow. We show that groups of swimmers enhance mixing, and puller-like swimmers are more efficient than pusher-like swimmers at olfactory shielding. Furthermore, we examine how swimmer kinematics and positional noise influence olfactory shielding. We find that puller-like swimmers remain more effective even under high positional noise. We also demonstrate that olfactory shielding is stronger for large swimmer Reynolds numbers and low flow Reynolds numbers, suggesting that olfactory shielding may contribute selective pressure on mesoscale swimmers in the ocean.
@@ II Modeling and methodology
We consider a collection of mesoscale swimmers, embedded in an open channel flow. The fluid flow is modeled through direct numerical simulation of the incompressible Navier-Stokes equation
@@
{align}
ρ(∂u∂t + u ⋅∇u) = -∇P + μ∇^2 u + f + f_p;
[…]
Figure 2: Statistics of the mean odor field. The contour plots show how the range of the mean odor field varies between pusher-like, puller-like and neutral swimmers for vertical and horizontal mid-plane cross-sections. (a) Vertical cross-sections of the mean odor differential $\Delta\bar{c}(\bm{x})$ for pusher-like swimmers and puller-like swimmers and the corresponding odor standard deviation differential $\Delta\sigma(\bm{x})$. The region containing the swimmers is marked by a red box. (b) The corresponding figures for the horizontal cross-section. (c) Odor profiles at different distances from the swimmers’ location. (d) Shielding intensities of the mean odor field $\Delta V(C)/V_{0}(C)$. (e) Shielding intensity of the odor fluctuations $\Delta V^{\sigma}(\theta)/V^{\sigma}_{0}(\theta)$ for pusher-like and puller-like swimmers, showing that the swimming dynamics dampens large fluctuations. (f) Impact of swimmer Reynolds number and (g) flow Reynolds number on the shielding intensity of the odor fluctuations $\Delta V^{\sigma}(\theta)/V^{\sigma}_{0}(\theta)$ for pusher-like and puller-like swimmers (evaluated at $\theta=0.1C_{\Omega}$).
The channel is forced with a constant pressure gradient $\bm{f}$ [28]. The boundary conditions for the channel flow are as follows: a fixed velocity on the bottom boundary (to represent the appropriate swimming velocity of the swimmers, as the simulations are carried out in the reference frame of the swimmers), free slip on the top boundary ($u_{z}=0,\partial_{z}u_{x}=\partial_{z}u_{y}=0$) and periodic in the other directions. We simulate three different fluid friction Reynolds numbers $Re_{\tau}=u_{\tau}H/\nu$, where $u_{\tau}$ is the shear velocity, $H$ is the channel height and $\nu$ is the kinematic viscosity ($\nu=\mu/\rho$), as shown in Table 1. We choose Environment 1 in Table 1 as our base environment and all units are normalized by using the Kolmogorov scales in this simulation. The simulation box has a dimension of $8000\eta\times 1000\eta\times 400\eta$, where $\eta$ is the Kolmogorov length scale.
@@
$\frac{\partial c}{\partial t}+\bm{u}\cdot\nabla c=D\nabla^{2}c+q.$ (1)
Here $c$ is odor concentration, $D$ is the diffusivity and $q$ is the odor source. We assume perfect adsorption on the bottom boundary ($c=0$), no flux at the top boundary ($\partial_{z}c=0$) and outflow conditions at the other boundaries. The source term $q$ at the position of the swimmers model a constant odor leakage rate. For our simulations, we use Schmidt number $Sc=1$. Although $Sc$ in water is typically higher than this, we anticipate a weak dependence on $Sc$ since we are interested in large-scale statistics [31, 33].
Figure 3: The effect of swimming dynamics on the detection probability. (a) $\Delta P(\bm{x};C_{H})$ for puller-like swimmers. (b) and (c) Odor pdf differential $\Delta f(c)$ in regions A and B, respectively, for puller-like swimmers. (d) Odor pdf differential $\Delta f(c)$ for both types of swimmers; statistics in regions A and B are represented with solid and dotted lines respectively. (e) Detection probability $P_{x}(x;C_{H})$ as a function of stream-wise distance, averaged along $y$ and $z$. The $x$-axis starts at $x=500$, where the group of swimmers ends. At long distances $x\gtrsim 700\,\eta$ the detection probability is reduced due to the swimming dynamics. Shielding persists up to $x\sim 2500\,\eta$.
[…]
To evaluate the effect of the kinematics of the swimmers on odor shielding, we conduct additional simulations where the swimmers are not fixed on a grid, but move according to the following Ornstein-Uhlenbeck process:
$d\bm{x}_{i}=-\frac{1}{\tau}(\bm{x}_{i}-\bm{\mu}_{i})dt+\sigma_{OU}\sqrt{dt}\bm{\eta_{i}},$ (2)
where $\bm{x}_{i}$ is the position of $i$th swimmer, $\bm{\mu}_{i}$ its equilibrium position, $\tau$ is the relaxation time, $\bm{\eta_{i}}$ the noise and $\sigma_{OU}$ is the strength of the noise. All three components of $\bm{\eta_{i}}$ are drawn from Gaussian distributions with zero mean and unit standard deviation. The equilibrium position of the swimmers are assumed to be the same as in the fixed grid configuration described above. However, unlike in the previous configuration, the swimmers can disperse, with $\sigma_{ms}=\sigma_{OU}\sqrt{\tau/2}$ defining the standard deviation of the swimmers’ position relative to their equilibrium positions.
@@
The list of environments is given in Table 1. All figures correspond to the base environment (Simulation 1 in Table 1) unless otherwise mentioned. For each of these environments, we conduct three distinct simulations: In the first two, we incorporate the dynamics of the swimmers by accounting for the forcing imposed by the swimmers on the fluid – pusher-like or puller-like; then we remove this forcing, providing a baseline against which to measure the swimmers’ effects. In the baseline simulation, the swimmers are retained as passive ‘ghosts’ that do not influence the flow, but emit odor at the same rate as the other two cases. The simulation setup is shown in Fig. 1, illustrating a typical case of odor dispersion in a channel flow.
@@ III Results
To determine whether swimmers’ dynamics leave their signatures in the odor distribution over long time-scales, we begin our analysis by examining the effect of hydrodynamic interactions on the mean and standard deviation of the odor field (Fig. 2). In this first part, we limit our analysis to swimmers fixed on a grid relative to the frame of reference, disregarding the effect of their kinematics. Since the odor distribution is inhomogeneous and statistically stationary, we approximate the ensemble average of the mean and standard deviation of the odor field with temporal averages: {align} ¯c(x) = ⟨c(x,t)⟩
σ(x) = ⟨(c(x,t) - ¯c(x,t))^2⟩^12. To account for the effect of the swimmers, we now define differential quantities by comparison with the baseline simulation that ignores the presence of the swimmers:
{align}
Δ¯c_s(x) = [¯c_s(x) - ¯c_0(x)]/C_Ω
Δσ_s(x) =[σ_s(x) - σ_0(x)]/C_Ω, where the subscript $s$ stands for swimmers (either pusher-like or puller-like), and the subscript 0 represents the baseline simulation. Mean and fluctuations are normalized by $C_{\Omega}=\text{max}(\bar{c}_{0})$, which occurs within the volume $\Omega$ containing the swimmers.
@@
Figure 5: Effect of swimmer kinematics on odor shielding under (a) intermediate positional noise (top row) and (b) high positional noise (bottom row), with snapshots of the swimmer positions is shown on the left.
To quantify the overall effect of swimmers in the entire volume, we define the volumes $V(C)$ where the mean odor concentration exceeds $C$ and $V^{\sigma}(\theta)$ where the fluctuations exceed $\theta$: {align} V_s(C) = ∫H(¯c_s(x) - C)dx,
V_s^σ(θ) = ∫H(σ_s(x) - θ)dx, where $H(x)$ is the Heaviside step function and integration is extended to the entire domain. We define the shielding efficiency as the decrease in the volumes defined above: {align} ΔV_s (C) = V_s (C) - V_0 (C),
ΔV_s^σ(θ) = V^σ_s (θ) - V^σ_0 (θ). When $C\approx 0$, swimming dynamics has little to no effect (or even a weak negative effect) in shielding the mean odor, with puller-like swimmers performing slightly worse than pusher-like ones (Fig. 2 (d)). As $C$ is increased, $\Delta V(C)<0$ thus both types of swimmers decrease the total volume where the average concentration is above the threshold. For a threshold of $C/C_{\Omega}=0.5$, pullers shrinks $V(C)$ by $50\%$, while pushers shrink $V(C)$ by $45\%$. Similarly, shielding efficiency for odor fluctuations (Fig. 2 (e)) shows that regions with extreme fluctuations are significantly reduced when the swimmer dynamics are taken into account. Interestingly, puller-like swimmers exhibit better shielding of both mean and fluctuations, which may reduce the chances of detection. These results stand in contrast to that of a single swimmer, which yields no significant shielding effect (Fig. S7).
Having examined odor shielding in the base environment, we now evaluate how the shielding changes as the Reynolds numbers of the swimmers and the advecting fluid are changed. At higher swimmer Reynolds numbers, the effect of forcing by the swimmers is expected to be more significant leading to an increase in shielding efficiency with $Re_{s}$. Our numerical experiments confirm this expectation (Fig. 2 (f) and Fig. S5). At lower swimmer $Re_{s}$ of 10 and 25, the swimmers’ dynamics have little effect on shielding the mean odor (Fig. S5 (a)). However, the fluctuations in the odor field do carry signatures of the swimming dynamics even at lower $Re$ for puller-like swimmers (Fig. 2 (f) and Fig. S5 (b)). To understand how the fluid $Re$ affects shielding, we repeat our analysis for two more fluid $Re$. Our analysis indicates that the effect of odor shielding is more prominent at lower fluid Reynolds numbers (Fig. 2 (g)). What is critical, however, is that puller-like swimmers continue to outperform pusher-like swimmers in all three fluid $Re$ considered in our analysis. A comparison of the swimmer and fluid Reynolds numbers that we consider with field observations and experiments of copepod swarms reported in Ref. [7] is provided in Table 2.
@@
Figure 6: Effect of swimmers’ kinematics on odor shielding when their mean position remains stationary relative to the ground. A snapshot of the swimmers’ positions is shown on the top right. Vertical and horizontal cross-sections of $\Delta\bar{c}(\bm{x})$ and $\Delta\sigma(\bm{x})$ for (a) pusher-like swimmers and (b) puller-like swimmers. (c) Shielding intensities of the mean odor field $\Delta V(C)/V_{0}(C)$. (d) Shielding intensity of the odor fluctuations $\Delta V^{\sigma}(\theta)/V^{\sigma}_{0}(\theta)$ for pusher-like and puller-like swimmers.
Unintuitively, the two swimmer types have opposite effects at the odor source, yet they both shrink their own detection probability at large enough distances from the group. This is best summarized by monitoring the probability of detection as a function of distance $x$ downstream of the swimmers: {align} P_x(x, C) = 1LzLy∫_0^L_z∫_0^L_y P(x, C)dydz. The group increases their probability of detection at close range, i.e. within $\sim$ $200\eta$ from the end of the group, about twice the length of the group. At longer range the group decrease their detection probability (Fig. 3 (e)). Importantly, shielding persists up to $\sim$ $2000\eta$, i.e. 20 times the length of the group. To reconcile this unintuitive mismatch, we analyze the full probability density function within regions A and B and compare the simulations with and without swimmers $\Delta f_{s}(c,\bm{x})=f_{s}(c,\bm{x})-f_{0}(c,\bm{x})$. In region A the swimmer dynamics dampen both extremes of the odor distribution (Fig. 3 (b)). This is a signature of increased mixing, which homogenizes the odor field towards its mean concentration. In region B the swimmer dynamics dilutes the odor, presumably by spreading it in a larger region further outward (Fig. 3 (c)). Pusher-like swimmers behave qualitatively similarly (Fig. 3 (d)). These signatures of enhanced mixing suggest that the velocity fluctuations due to the swimmer dynamics play a major role in odor transport. The effect of swimmer and flow Reynolds numbers on the probability of detection is shown in Fig. S5. As we already observed for the mean odor field, shielding of the probability of detection is more effective at higher swimmer Reynolds numbers and lower flow Reynolds numbers.
@@
Study Swimmer Reynolds number $Re_{s}$ Bulk Reynolds number of the flow $Re_{b}$
Present work 10 – 50 6,000 – 18,050
Laboratory experiment in Ref. [7] $\leq$20 $\leq$2,000
Field observations in Ref. [7] $\leq$20 $\leq$15,000
Table 2: Comparison of the swimmer Reynolds number $Re_{s}$ and the bulk Reynolds number $Re_{b}=U_{b}H/\nu$, where $U_{b}$ is the mean flow velocity and $H$ is the flow depth, between the present study and the experimental and field observations of copepod swarms reported in Ref. [7].